OpenStudy (andriod09):

(x-4)(x-(1/4))(x+10)=(log(pi)+sin(pi/7)^0-1

OpenStudy (dape):

So it looks pretty complicated, but if you take a look at the right side you see that the complex pi/log/sin stuff is all raised to the 0th power, which means it's 1 (ANYTHING raised to the 0th power is 1). So if we replace that by 1 we get \[ (x-4)(x-\frac{1}{4})(x+10)=1-1=0 \] Now it should be pretty easy.

OpenStudy (andriod09):

I CANNOT SEE EQUATIONS FRO MTHE EQUATION BUTTON!!! ALL I SEE ARE THE SLASHES AND THE BRACKETS AND THE BRACES!!

OpenStudy (dape):

Okay, what I wrote was: (x-4)(x-1/4)(x+10)=1-1=0

OpenStudy (andriod09):

and also, i kthe answer, i was wondering how many others can get it.